Farmers keeps 1000 Jersey cows and each produces 20 litres of milk that is sold at R5.00 per litre.About 100 litres consumed at the farmer’s home while 1000 litres were donated to orphanage. About 30 cows were sold at M5000.00 each while 15 cows were donated to the farm at M4950.00 for each cow. About 200 litres were stolen by two notorious farm workers.


About 10 000kg of grain was bought from farm at M500.00 per tonne and used for supplementary feeding. Milkers under the supervision of chief milker charged the rate of M2.00 per cow. Veterinary nurse charged M3000.00 for her services and drugs. About 5000 packaging bottles of 5 litre size were bought and NOT ALL were used and each cost M2.00. The marketing costs were 10% of the value of milks sales and 5% of the value of cow sales.

From the above information;

1. Calculate the gross income for the enterprise.

2. Calculate gross margin for the enterprise.


1
Expert's answer
2021-05-13T17:45:45-0400

1 : Gross Income for the enterprise = Sales Revenue - Cost of Goods Sold. 

1 tonne = 1000 kg 

10,000 kg = 10 tonne 

Cost of 10,000 kg of grain "= 10 \times 500 = M5,000"

Veterinary nurse "= M3,000"

Number of Cows remaining for milk extraction : 

"= 1,000 - 15 - 30"

"= 955\space cows"

Milker charges "= 955 \times 2 = M1,910"

Milk produced "= 955 \times 20 = 19,100 \space litres"

Milk sold "= 19,100 - 100\space (consumed \space at\space home) - 1,000\space (donated) - 200 \space (stolen)"

"= 17,800 \space litres"

Bottles to be used for packaging "= \frac{17,800 }{ 5} = 3,560"

Cost of packaging "= 3,560 \times M2 = M7,120"

Milk sales revenue "= 17,800 \times M5 = M89,000"

Cost of Milk sold = Cost of grains + Veterinary Nurses charges + Milker Charges + Packaging charges. 

"= M5,000 + M3,000 + M1,910 + M7,120"

"= M17,030"

Gross Income for the enterprise = Sales Revenue - Cost of Goods Sold. 

"= M89,000 - M17,030"

"= M71,970"

 

2 : Gross Margin for the enterprise :

"= \frac{Gross \space Income }{Net\space sales}\times 100"


"= \frac{M71,970 } {M89,000} \times100"


"= 80.86\%"